OpenStudy (anonymous):

A funnel with a height of 3.67 inches is shown below. The funnel has a diameter of 1.87 inches that includes a wall that is 0.125 inches thick. What is the maximum volume of the funnel rounded to the nearest hundredth? Answer 7.57 in3 3.36 in3 2.52 in3 2.22 in3

OpenStudy (swag):

b

OpenStudy (anonymous):

are you serious? or are you guessing?

OpenStudy (swag):

i neve rguess

OpenStudy (anonymous):

Hello? castillo you still there?

OpenStudy (callisto):

yes... just fed up with the calculation :( gimme some time

OpenStudy (anonymous):

sprry! just making sure

OpenStudy (callisto):

i dunno .. my answer isn't even in one of these options :(

OpenStudy (callisto):

so the formula is V=(1/3)π(r^2)h ?

OpenStudy (anonymous):

Yes. I am having trouble as well. Im thinking I need to find the volume of both and add them together. the inner and outer

OpenStudy (callisto):

nah... that's not possible..

OpenStudy (callisto):

the vol. of cone without the wall = (1/3)π(r^2)h = (1/3)π[(1.87-0.125x2)/2]^2 (3.67) = (1/3)π (0.81)^2 (3.67) = 2.52

OpenStudy (anonymous):

MEDALS?

OpenStudy (anonymous):

I thought it was bs also

OpenStudy (anonymous):

sorry "b"

OpenStudy (anonymous):

For futur ref please put the options synchronized properly. Answer 7.57 in3 3.36 in3 2.52 in3 2.22 in3 to: Answer 7.57 in3 3.36 in3 2.52 in3 2.22 in3

OpenStudy (anonymous):

but the question says the wall is included in the diameter

OpenStudy (anonymous):

Ohh ok sory haha! thats how it comes out sometimes

OpenStudy (anonymous):

Duh Duh, I am sorry, Castillo is right.

OpenStudy (anonymous):

\[\frac 13 \times \frac {(1.87-2 \times 0.125)}{2}^2 \times \pi \times 3.67 = 2.52153 \]

OpenStudy (anonymous):

i got ya fool foolformath

OpenStudy (anonymous):

you gys are smart in here. you must have like A's all throughout geometry and chemistry

OpenStudy (callisto):

poor me.. no one understand what i wrote :(

OpenStudy (anonymous):

no I do.what is the volume of a semicircle